# 1.1 Preliminaries on probes

### v1

@@ 1 Quantum and Classical Probes over a Channel
In this section, we consider quantum and classical probes that are sent through a lossy channel with transmissivity $\eta_{0}\in(0,1)$. Suppose at an unknown time $\nu$, the transmissivity of the channel drops from $\eta_{0}=\eta\in(0,1)$ to $\eta_{1}=\eta_{0}\eta_{d}$, $\eta_{d}\in(0,1)$. We first review existing results from [guha2025quantum] in § 1.1, and then proceed to derive the quantum advantage in reducing detection delay over a channel in § 1.2.
@@ 1.1 Preliminaries on probes
We begin with a brief introduction to the quantum and classical probes used in this paper, followed by a short description of how the probes are used to detect transmissivity change in a channel.
@@
As in [guha2025quantum], the classical (or quasi-classical) benchmark we consider is a stream of $T$ coherent-state pulses |$\ket{\alpha}^{\otimes T}$, with $\alpha\in\mathbb{R}$ and mean photon number $\alpha^{2}=N+N_{a}$. Here $0<N_{a}<<N$ is introduced to fairly compare the classical case with its quantum-augmented counterparts. At the receiving monitor, |$\ket{\alpha}^{\otimes T}$ generates $T$ i.i.d. Gaussian-distributed random variables $X_{1},X_{2},\dots,X_{T}$, where
@@
Next, we consider the entanglement-augmented probes. Specifically, $T$ blocks of $n\in\mathbb{Z}_{+}$ coherent-state pulses |$\ket{\alpha}^{\otimes n}$ are augmented by a continuous variable entangled state generated by splitting a squeezed-vaccum state |$\ket{0;s}$, $s\in\mathbb{R}_{+}$, where
@@
In particular, when $n=1$, we have a stream of $T$ displaced squeezed states |$\ket{\alpha;r}^{\otimes T}$, with $\alpha^{2}=N$ and $\sinh^{2}(r)=N_{a}$. The receiving monitor outputs $T$ i.i.d. Gaussian-distributed random variables $X_{1},X_{2},\dots,X_{T}$, where
@@
and $\sigma_{j}^{2}=\frac{1}{4}(\eta_{j}e^{-2r}+1-\eta_{j})$, and $j=0$ pre-change, $j=1$ post-change. While the physical processes of generating the probes for $n=1$ and $n\geq 2$ are different [guha2025quantum], mathematically, there is in fact no need to distinguish these two cases. We only make the distinction when discussing previous results in relation to ours in § 1.2. Whenever necessary, we refer to $n=1$ as the squeezing augmentation case, and $n\geq 2$ the general quantum augmentation case.
@@
$\text{CUSUM}[t]=\max\left\{0,\text{CUSUM}[t]+\log{\frac{f_{1}(X_{t})}{f_{0}(X_{t})}}\right\},\forall t\in[T],$
and $\text{CUSUM}[0]=0$. Let $\gamma$ be the minimum time to false alarm that one is willing to tolerate, and denote $\tau$ as the minimum $t$ such that $\text{CUSUM}[t]>\log{\gamma}$. Then the CUSUM algorithm outputs $\tau$ as the estimate for the change point $\nu$. Informally, as $\gamma\to\infty$, the CUSUM algorithm achieves a worst-case detection delay [L1971] of
@@
$D\left(\mathcal{N}_{n}^{(1)}||\mathcal{N}_{n}^{(0)}\right)=\frac{1}{2}\left(\operatorname{Tr}\left({\left(\bm{\Sigma}^{(0)}\right)^{-1}\bm{\Sigma}^{(1)}}\right)+\left(\bm{\mu}^{(0)}-\bm{\mu}^{(1)}\right)^{T}\left(\bm{\Sigma}^{(0)}\right)^{-1}\left(\bm{\mu}^{(0)}-\bm{\mu}^{(1)}\right)-n+\ln{\frac{\left\lvert{\bm{\Sigma}^{(0)}}\right\rvert}{\left\lvert{\bm{\Sigma}^{(1)}}\right\rvert}}\right),$ (8)
@@
1.2 The quantum advantage in reducing detection delay
@@
First note that from (5), covariance matrix $\bm{\Sigma}^{(j)}$ has uniform diagonal entries $d_{j}=\frac{1}{4}\left(1-\frac{\eta_{j}(1-e^{-2s})}{4n}\right)$ and uniform off-diagonal entries $r_{j}=d_{j}-\frac{1}{4}$. Such a structure allows us to explicitly write out both $\left\lvert{\bm{\Sigma}^{(j)}}\right\rvert$ and $\left(\bm{\Sigma}^{(j)}\right)^{-1}$.
@@
Using the explicit form of the per-pulse quantum KL divergence (16), we are able to rigorously derive several results that have been previously inferred through numerical evaluations in [guha2025quantum]. It has been shown that the KL divergence sees a sharp increase when augmenting a tiny amount of squeezing. That is, when $n=1$, $\lim_{N_{a}\to 0}\frac{\partial}{\partial N_{a}}\frac{1}{n}D\left(\mathcal{N}_{n}^{(1)}||\mathcal{N}_{n}^{(0)}\right)=\lim_{N_{a}\to 0}\frac{\partial}{\partial N_{a}}D\left(\mathcal{N}_{1}^{(1)}||\mathcal{N}_{1}^{(0)}\right)=\infty$. It has also been observed without proof that the same holds for general $n$. By taking the partial derivative of (16), it is straightforward to verify that indeed,
@@
Additionally, numerical evaluations (e.g. Figure 2 in [guha2025quantum], with $n=2^{i}$, $i\in[8]$) suggest that for fixed $N$, $N_{a}$, $\eta$ and $\eta_{d}$, the quantum KL divergence $\frac{1}{n}D\left(\mathcal{N}_{n}^{(1)}||\mathcal{N}_{n}^{(0)}\right)$ converges to $\frac{2N\eta(1-\sqrt{\eta_{d}})^{2}}{1-\eta}$ as $n$ becomes large. This directly follows from taking the limit of (16) as $n$ goes to infinity.
With Lemma 1 in hand, we are well-equipped to understand, given a set of parameters, whether quantum probes outperform the classical ones, and if so, by how much. Moreover, we can explore how the quantum advantage changes with respect to each variable, and what is the asymptotic behavior of such advantage.

### v2

# A Quantum Speedup in Localizing Transmission Loss Change in Optical Networks

@@
Abstract
The ability to localize transmission loss change to a subset of links in optical networks is crucial for maintaining network reliability, performance and security. Quantum probes, implemented by sending blocks of $n$ coherent-state pulses augmented with continuous-variable (CV) squeezing ($n=1$) or weak temporal-mode entanglement ($n>1$) over a lossy channel to a receiver with homodyne detection capabilities, are known to be more sensitive than their quasi-classical counterparts in detecting a sudden increase in channel loss. The enhanced sensitivity can be characterized by the increased Kullback-Leibler (KL) divergence of the homodyne output, before and after the loss change occurs. When combined with the theory of quickest change detection (QCD), the increase in KL divergence translates into a decrease in detection latency.
In this work, we first revisit quantum probes over a channel, and consider the ratio of quantum to classical KL divergences as a metric for quantum speedup. We generalize previous results on $n=1$ (CV squeezed states) to arbitrary values of $n$, and discuss the monotonicity and asymptotics of this ratio with respect to the parameters of the quantum probe. Assuming a subset of nodes in an optical network is capable of sending and receiving such probes through intermediate nodes with all-optical switching capabilities, we present a scheme for quickly detecting the links that have suffered a sudden drop in transmissivity. Since quantum probes lose their sensitivity with increasing loss in the channel, we first propose a probe construction algorithm that makes the set of links suffering transmission loss change identifiable, while minimizing the longest distance a probe traverses. We then introduce new cumulative sum (CUSUM) statistics with a stopping rule, which allows us to run the CUSUM algorithm to quickly localize the faulty links using our constructed probes. Finally, we show that the proposed scheme achieves a quantum speedup in decreasing the localization delay.
1 Introduction
If we augment part of a classical network with quantum capabilities, what can we do better than in the purely classical setting? In this paper, we show that detecting transmission loss change in an optical network provides one answer to this question.
Optical networks are highly prevalent in modern communication systems. Major cloud providers rely on optical networks to connect data centers and ensure reliable access to cloud services [17, 27, 33]. Additionally, optical mesh networks, e.g., Reconfigurable Optical Add-Drop Multiplexer (ROADM) networks have seen significant density of deployment by metropolitan area Internet Service Providers and telecom providers. However, unattended optical fiber networks are often subjected to physical damages to fibers, or unauthorized fiber tapping [23, 30, 16]. These lead to increases in transmission loss, which often translates into increased bit error rates, which may in turn impact the quality of service. Perhaps more strikingly, by introducing only a small transmission loss, an eavesdropper can extract information on the data being transferred [23, 16]. Therefore, the ability to detect transmission loss in an optical network is crucial for maintaining network reliability, performance and security.
Classical and quasi-classical methods such as bit error rate measurement, optical spectrum analysis (OSA), and mean optical power monitoring exist for detecting changes in transmission loss in an optical channel, but are generally not fast enough to detect a change in transmission loss that lasts for only seconds [30, 11]. Optical time-domain reflectometry (OTDR) is used for fault detection and localization, but it is not sensitive enough for the order of loss incurred by a stealthy eavesdropper [3, 29, 11]. The Quantum Alarm technique [11] achieves better empirical sensitivity, by using the parameter estimation step of continuous-variable quantum key distribution (CV-QKD) protocols to act as an alarm for loss detection. The quasi-classical benchmark in [12] is a better protocol due to the data processing inequality—while the parameters used in Quantum Alarm [11] are the output of homodyne detection after post-processing—the quasi-classical benchmark in [12] works directly with the probabilistic distributions from the homodyne output, using the theory of quickest change detection (QCD).
On the quantum side, [12] recently shows that, a vanishingly small quantum energy in the form of squeezed states, when added to the bright classical pulses, dramatically increases the sensitivity in detecting transmission loss change in an optical channel. The quantum protocols in [12] are simple to implement and effective, with the simplest protocol involving sending displaced squeezed states as probes from a squeezing augmented transmitter over a lossy channel, and processing the output at a homodyne detection receiver. In the advanced protocol, coherent states augmented by a continuous-variable entanglement produced by a temporal-mode Green Machine [7] are sent as probes, and the outputs are similarly captured by a homodyne detector. Feeding the homodyne outputs into a cumulative sum (CUSUM) algorithm [25], one can quickly catch the change, with a detection delay that is inversely proportional to the Kullback-Leibler (KL) divergence from the pre-change to the post-change distributions. Compared to the quasi-classical benchmark, the quantum probes induce a larger KL divergence, which leads to a lower detection delay.
In this paper, we first revisit quantum probes over a channel, generalize and extend existing theoretical results on squeezing-augmented probes to the general entanglement-augmented probes. Then, building on the quantum probes for a channel [12], we propose a method for detecting and localizing transmission loss change in a network. Though the technologies for sending and receiving these probes are readily available, it is only realistic to assume that a subset of the network nodes are equipped with such quantum capabilities. We refer to the quantum-capable nodes as monitors. We assume the rest of the nodes, the non-monitors, are passive optical switches that can pass the quantum probes through.
When expanding the goal from change detection in a channel to fault localization in a network, we must decide how to route the probes given a network topology and a set of monitors. If we think of whether a link suffers a sudden transmission drop as a Boolean state of the link, the framework of Boolean network tomography [8] can help infer every link’s state with limited visibility through probes between monitors, and therefore localize the transmission drop. While the probes do not output Boolean values, their ability to distinguish whether a change occurs on the links they traverse is what makes the Boolean framework relevant. Previous works have studied where to place monitors in a network [21], conditions that allow network failures to be identified [22], and how to construct probes to localize faults [1, 32]. Recognizing that we often do not get to choose where quantum capabilities are placed, in this work, we focus on the probe construction problem. The fact that probes are quantum brings in new objectives in designing the construction algorithms. Due to factors such as attenuation, noise and decoherence, quantum probes are inherently sensitive to the the distance they travel. While existing heuristics aim to minimize the total cost [1, 32] or the total length of probes in the context of classical networks, we introduce a new probe construction algorithm that optimizes the distance the longest probes traverse.
Assuming all observations from the probes are collected at a central site, we develop a network-level CUSUM algorithm to quickly identify the problematic links. Despite the CUSUM algorithm being a typical choice for detecting change [12, 36, 35], much less is known on how to apply it in the setting of fault localization in a network. For simplicity, we restrict our attention to the case where at most one link in a network suffers from a transmission drop. To the best of our knowledge, the only previous attempt is a CUSUM-based approach proposed to identify outages in a power network [28], which outputs candidates of problematic links, and relies on further simulations to narrow down the search. Using the probes generated by our construction algorithm, we are able to adapt the CUSUM algorithm to directly pinpoint the faulty link.
Our contributions in this paper are twofold:
  • We generalize and extend existing results on squeezing-augmented probes to the general entanglement-augmented probes, and characterize the quantum speedup in reducing change detection delay over a channel (§ 2).
  • We present a scheme with quantum speedup in localizing transmission drop in a network, which incorporates a probe construction algorithm (§ 3), and a CUSUM-based fault localization algorithm (§ 4).
We evaluate the performance of our network-level CUSUM algorithm on both a line network, and an all-optical version of the 3-tier fat-tree network topology [2] (§ 5). Finally, we discuss related work in § 6, and conclude the paper in § 7.
@@ 2 Quantum and Classical Probes over a Channel
In this section, we consider quantum and classical probes that are sent through a lossy channel with transmissivity $\eta_{0}\in(0,1)$. Suppose at an unknown time $\nu$, the transmissivity of the channel drops from $\eta_{0}=\eta\in(0,1)$ to $\eta_{1}=\eta_{0}\eta_{d}$, $\eta_{d}\in(0,1)$. We first review existing results from [12] in § 2.1, and then proceed to derive the quantum speedup in reducing detection delay over a channel in § 2.2.
@@ 2.1 Preliminaries on probes
We begin with a brief introduction to the quantum and classical probes used in this paper, followed by a short description of how the probes are used to detect a transmissivity drop in a channel.
@@
[…]
@@
[…]
@@
[…]
@@
and $\sigma_{j}^{2}=\frac{1}{4}(\eta_{j}e^{-2r}+1-\eta_{j})$, and $j=0$ pre-change, $j=1$ post-change. While the physical processes of generating the probes for $n=1$ and $n\geq 2$ are different [12], mathematically, there is in fact no need to distinguish these two cases. We only make the distinction when discussing previous results in relation to ours in § 2.2. Whenever necessary, we refer to $n=1$ as the squeezing augmentation case, and $n\geq 2$ the general quantum augmentation case.
@@
$\text{CUSUM}[t]=\max\left\{0,\text{CUSUM}[t-1]+\log{\frac{f_{1}(X_{t})}{f_{0}(X_{t})}}\right\},\forall t\in[T],$
and $\text{CUSUM}[0]=0$. Let $\gamma$ be the minimum time to false alarm that one is willing to tolerate, and denote $\tau$ as the minimum $t$ such that $\text{CUSUM}[t]\geq\log{\gamma}$. Then the CUSUM algorithm outputs $\tau$ as the estimate for the change point $\nu$. Informally, as $\gamma\to\infty$, the CUSUM algorithm achieves a worst-case detection delay [20] of
@@
[…]
@@
2.2 The quantum speedup in reducing detection delay
@@
First note that from (5), covariance matrix $\bm{\Sigma}^{(j)}$ has uniform diagonal entries $d_{j}=\frac{1}{4}\left(1-\frac{\eta_{j}(1-e^{-2s})}{4n}\right)$ and uniform off-diagonal entries $r_{j}=d_{j}-\frac{1}{4}$. Such a structure allows us to explicitly write out both $\left\lvert{\bm{\Sigma}^{(j)}}\right\rvert$ and $\left(\bm{\Sigma}^{(j)}\right)^{-1}$, where
@@
Using the explicit form of the per-pulse quantum KL divergence (16), we are able to rigorously derive several results that have been previously inferred through numerical evaluations in [12]. It has been shown that the KL divergence sees a sharp increase when augmenting a tiny amount of squeezing. That is, when $n=1$, $\lim_{N_{a}\to 0}\frac{\partial}{\partial N_{a}}\frac{1}{n}D\left(\mathcal{N}_{n}^{(1)}||\mathcal{N}_{n}^{(0)}\right)=\lim_{N_{a}\to 0}\frac{\partial}{\partial N_{a}}D\left(\mathcal{N}_{1}^{(1)}||\mathcal{N}_{1}^{(0)}\right)=\infty$. It has also been observed without proof that the same holds for general $n$. By taking the partial derivative of (16), it is straightforward to verify that indeed,
@@
Additionally, numerical evaluations (e.g. Figure 2 in [12], with $n=2^{i}$, $i\in[8]$) suggest that for fixed $N$, $N_{a}$, $\eta$ and $\eta_{d}$, the quantum KL divergence $\frac{1}{n}D\left(\mathcal{N}_{n}^{(1)}||\mathcal{N}_{n}^{(0)}\right)$ converges to $\frac{2N\eta(1-\sqrt{\eta_{d}})^{2}}{1-\eta}$ as $n$ becomes large. This directly follows from taking the limit of (16) as $n$ goes to infinity.
With Lemma 1 in hand, we are well-equipped to understand, given a set of parameters, whether quantum probes outperform the classical ones, and if so, by how much. Moreover, we can explore how the quantum speedup factor changes with respect to each variable, and what is the asymptotic behavior of such speedup.
@@
      $\begin{aligned} N_{a,0}=&\frac{8Nn\left(3\eta-1\right)+b_{d}\left(4\eta-1\right)}{4n\left(8Nn\left(1-\eta\right)-\eta b_{d}\left(1-\eta_{d}\right)\right)}\\ &+\frac{\sqrt{8Nn\left(b_{d}+4Nn\eta\right)\left(8Nn\left(1-\eta\right)-\eta b_{d}\left(1-\eta_{d}\right)\right)+\left(8Nn\left(3\eta-1\right)+b_{d}\left(4\eta-1\right)+4\right)^{2}}}{4n\left(8Nn\left(1-\eta\right)-\eta b_{d}\left(1-\eta_{d}\right)\right)}.\end{aligned}$
For all claims on the monotonicity, we bound the corresponding partial derivatives. We refer the reader to Appendix A in the full version of our paper [34] for the proof of Lemma 2.
Lemma 2(1)(3) not only corroborates an observation that quantum augmentation provides the most benefit in the large-$N$ and high-$\eta$ regime [12] when $n=1$, but also shows that the same holds in the general quantum augmentation case (arbitrary $n$). Additionally, [12] provides an approximate threshold $\frac{N\eta}{1-\eta}$ on $N_{a}$, above which squeezing augmentation does not outperform the classical benchmark in the case of $n=1$. Lemma 2(5)(5.1) gives a range of $N_{a}$ where quantum speedup exists, for any $n$. We note that, for typical values of $N$, $n$, $\eta$ and $\eta_{d}$ (e.g., see § 5 and [12]), $q_{n}$ is likely to first increase in $N_{a}$, and then decrease. Lacking ways to directly find the $N_{a}$ that maximizes $q_{n}$, we settle for the result in Lemma 2(5)(5.3). Nonetheless, $N_{a,0}$ may be of practical interest, as it narrows down the range of $N_{a}$ in which the optimal quantum speedup factor appears. Too see that, given typical values such as $N=100$, $n=1$, $\eta=0.9$ and $\eta_{d}=0.8$, Lemma 2(5)(5.1) gives that the quantum speedup exists when $0<N_{a}<897.5$ by taking $\epsilon$ ranging from arbitrarily close to 0 to $890$, and [12] shows that such speedup disappears when $N_{a}$ is above roughly $900$; however, both bounds on $N_{a}$ are rendered irrelevant under the assumption that $N_{a}<<N=100$. In contrast, Lemma 2(5)(5.3) indicates that quantum augmentation brings the most value when $N_{a}<N_{a,0}=20.9$. This is evidence that assuming $N_{a}<<N$ (§ 2.1) makes sense.
Beyond generalizing and strengthening previous results, we can draw a few interesting conclusions from Lemma 2. While smaller transmissivity drops (larger $\eta_{d}$) are generally harder to detect, the quantum speedup factor increases as the drop gets lighter (Lemma 2(2)). Moreover, with proper choices of parameters, the quantum speedup exists, and the limits in Lemma 2(2)(3)(4) provide the respective upper bounds on how much quantum augmentation could help.
3 Probe Construction for a Network
A network is a weighted graph $G=(V,E)$, with a set of monitors $M\subseteq V$ that can send and receive probes. If we think of each edge $e$ in $G$ as a channel with transmissivity $\eta_{e}$, the edge weight $w(e)$ of $e$ is set to $\log{\eta_{e}}$. Throughout this paper, we use the term link in the context of a network, but refer to a link as an edge when we discuss the network in graph-theoretic terms. In this section, we first formulate the problem in the language of Boolean network tomography (§ 3.1), and later introduce an algorithm for constructing the probes (§ 3.2).
3.1 The probe construction problem
Throughout this section, we refer to transmissivity drops as faults, and assume only a subset of $E$ in $G$ may be faulty. A probe $P$ is a walk [^1] A walk with endpoints $v_{0}$ and $v_{k}$ is a sequence $v_{0},e_{1},v_{1},e_{2},...,e_{k},v_{k}$, where $v_{i}\in V$ for all $0\leq i\leq k$, and $e_{j}\in E$ for all $1\leq i\leq k$. with endpoints in the set of monitors MM. [^2] This definition corresponds to the Arbitrarily Controllable Routing (ACR) scenario in the classical tomography literature (§ 1.4 in [15]). For example, cycles are allowed, and each edge can be traversed in both directions. When the context is clear, we also think of $P$ as the set of edges the probe passes through, and the length of $P$ is $l(P)=\sum_{e\in P}w(e)=-\log{(\prod_{e\in P}\eta_{e})}$. A probe $P$ is faulty, if an only if there exists an edge $e\in P$ that is faulty. As we shall see in § 4, the state of a probe is in fact an observation drawn from the pre- and post-change distributions described in § 2. And the Boolean nature of the probes described in this section corresponds to whether the pre- and post-change distributions differ.
We call $F\subseteq E$ a fault set, if all edges in $F$ are simultaneously faulty. For a set of probes $\mathcal{P}$, the fault set $F$ is manifested through $\mathcal{P}_{F}=\{P\in\mathcal{P}\mid P\cap F\neq\emptyset\}$. Two fault sets $F_{1}$ and $F_{2}$ are distinguishable if and only if $\mathcal{P}_{F_{1}}\neq\mathcal{P}_{F_{2}}$. A graph $G$ is identifiable with respect to a set of all possible faults $\mathcal{F}$ given a set of monitors $M$, if for any pair of distinct fault sets $F_{1},F_{2}\in\mathcal{F}$, $F_{1}$ and $F_{2}$ are distinguishable. In practice, we often do not get to select the network topology (the graph $G$), which nodes are quantum capable and can serve as monitors (monitor set $M$), and what part of a network may be faulty (the set of all possible faults $\mathcal{F}$). For simplicity, we assume $G$ is identifiable w.r.t. the given $M$ and $\mathcal{F}$. [^3] Relaxing this assumption is straightforward, the proposed algorithm can be easily adapted to identify a maximal subset of $\mathcal{F}$. And our goal is to construct a set of probes $\mathcal{P}$, such that every pair of distinct fault sets in $\mathcal{F}$ are distinguishable.
The problem of constructing $\mathcal{P}$ is very well-studied, both in the context of classical network tomography [5, 32], and combinatorial group testing [4, 14, 31], with the former literature focusing on finding heuristics to minimize the total length $\sum_{P\in\mathcal{P}}l(P)$ of the probes, and the latter putting an emphasis on the optimality of $\left\lvert{\mathcal{P}}\right\rvert$. The challenge posed by our problem is that, quantum probes are inherently sensitive to the distance they travel. As a probe traverses through multiple edges in the graph, the transmissivities of edges (channels) multiply, and from § 2.2, it is always desirable to avoid long probes and low transmissivity whenever possible. Therefore, we are interested in finding the set of probes $\mathcal{P}^{*}$ that minimizes the length of the longest probe:
Problem 1 (Probe construction).
Construct $\mathcal{P}^{*}=\arg\min_{\mathcal{P}}\max_{P\in\mathcal{P}}l(P)$, where $\mathcal{P}$ is a set of probes such that every pair of distinct fault sets in $\mathcal{F}$ are distinguishable given $G$ and $M$.
3.2 Optimizing the longest probes
The meta algorithm.
Existing algorithms for constructing $\mathcal{P}$ commonly follow a meta algorithm shown as Algorithm 1 [15, Algorithm 25]. This meta algorithm iteratively goes through each pair of distinct fault sets $F_{1}$ and $F_{2}$, and invokes a subroutine (Step 5) to add a probe that distinguishes $F_{1}$ and $F_{2}$. Here the tag $t_{F}$ (Step 2) keeps track of $\mathcal{P}_{F}$ as the algorithm proceeds, and can be thought of as the syndrome of the fault set $F$. If $t_{F_{1}}\neq t_{F_{2}}$, $F_{1}$ and $F_{2}$ are distinguishable, as $\mathcal{P}_{F_{1}}\neq\mathcal{P}_{F_{2}}$. Also observe that once $t_{F_{1}}\neq t_{F_{2}}$ starting from some iteration $i$, they remain unequal till the end, as $t_{F_{1}}$ and $t_{F_{2}}$ differ in the $i$ less significant bits.
Algorithm 1 A meta algorithm for probe construction [15]
Input: A graph $G=(V,E)$, a set of all possible faults $\mathcal{F}$, and a set of monitors $M$;
Output: A set of probes $\mathcal{P}$.